16: What is the volume of a right cylinder with radius 4 cm and height 3 cm? (BLANK) (pi) cm3
Answer by dolly(163)

Volume of the cylinder =Pi (r)^2h
= (22/7)4^2*3
= (22/7)(16)*3
= 150.84 cu. cm